give the dimesions of five rectangles tha have a perimeter of 50
Answer by stanbon(75887)

give the dimesions of five rectangles tha have a perimeter of 50
---------------
Perimeter = 2(L+W)
50 = 2(L+W)
L+W = 25
--------
L=1,W=24
L=2,W=23
L=3,W=22
L=5,W=20
etc.
=====================
Cheers,
Stan H.